Ask
your own question, for FREE!
Mathematics
9 Online
two cards are drawn without replacement from a standard 52 card deck. what is the probability of drawing two kings?
Still Need Help?
Join the QuestionCove community and study together with friends!
i have 1/169 ; 1/221 ; 4/663 ; 3/676 as a choice of an answer
The probability of drawing the first card is 4/52. Since the card has not been replaced, the probability of drawing the second card is 3/51. Thus the probability of drawing 2 kings without replacement is (4/52)(3/51) =1/221 = .00452.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
3 days ago
3 Replies
2 Medals
6 days ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal